OpenStudy (sabrinaaurora):

Brenda wants to make 160 photocopies that each cost $0.045 WHAT IS her total before tax

OpenStudy (christopher1448):

52cents

OpenStudy (retireed):

A hundred copies would be $4.50. So 52 cents is a little lite.

OpenStudy (anthonyym):

1 photocopy = 0.045 2 photocopies = 0.045 + 0.045 = 2(0.045) 3 photocopies = 0.045 + 0.045 + 0.045 = 3(0.045) Cost for n photocopies = n(0.045)

OpenStudy (cayla):

$7.20

OpenStudy (bonnieisflash1.0):

scdhaSB

OpenStudy (joshmorris):

lethbw